# Gravitational collapse of a degenerate wormhole

### v1

@@ 8 Conclusions
Due to spherical symmetry, the solution to this problem is significantly simplified by the possibility of applying the equivalence principle - extended to account for the presence of matter-free sources of the gravitational field. Direct application of the extended equivalence principle reduces the radial collapse problem of the Klinkhamer wormhole to the well-known problem of a test particle falling in a Schwarzschild gravitational field. This uniquely determines the dynamical equation for $b(t)$. As follows from the phase portrait in Fig. 2, the corresponding Cauchy problem has a unique solution for any initial conditions ($b_{0},\dot{b_{0}}$).

### v2

@@ 8 Conclusions
Due to spherical symmetry, the solution to this problem is significantly simplified by the possibility of applying the equivalence principle - extended to account for the presence of matter-free sources of the gravitational field. Direct application of the equivalence principle reduces the radial collapse problem of the Klinkhamer wormhole to the well-known problem of a test particle falling in a Schwarzschild gravitational field. This uniquely determines the dynamical equation for $b(t)$. As follows from the phase portrait in Fig. 2, the corresponding Cauchy problem has a unique solution for any initial conditions ($b_{0},\dot{b_{0}}$).